OpenStudy (anonymous):

Simplify 3/x-1 - 2/x

OpenStudy (anonymous):

is it \[x+2/x^2 - x \] ??

hero (hero):

As you posted it, your expression is \[\frac{3}{x} - 1 - \frac{2}{x}\] But you may have meant for it to be \[\frac{3}{x - 1} - \frac{2}{x}\]

hero (hero):

Also your result, as you posted it is \[x + \frac{2}{x^2} - x\] But perhaps you meant to say \[\frac{x - 2}{x^2 - x}\]

OpenStudy (anonymous):

\[\frac{ 3 }{ x-1 } - \frac{ 2 }{ x }\] yes that's what I meant, and the answer,

hero (hero):

Multiply the first fraction by \(\dfrac{x}{x}\) Multiply the second fraction by \(\dfrac{x - 1}{x - 1}\) to get \(\dfrac{3x}{x(x - 1)}-\dfrac{2(x - 1)}{x(x - 1)} = \dfrac{3x - 2(x - 1)}{x(x - 1)} = \dfrac{3x - 2x + 1}{x(x - 1)} = \dfrac{x + 1}{x(x - 1)}\)

OpenStudy (anonymous):

but -2(x-1) = -2x +2 ? so I cant see how you get x +1 for the numerator ? geez im gonna look silly :)

OpenStudy (anonymous):

should be 3x - 2x + 2 in the numerator !!!

hero (hero):

Yes, you are right

hero (hero):

So \(\dfrac{x + 2}{x(x - 1)}\)

OpenStudy (anonymous):

so if you look at my original answer it ok, :) thx for help

hero (hero):

Yes, I know. I was just trying to type out the steps but I made a typo. Typed 1 instead of 2.

OpenStudy (anonymous):

yes, i gathered that, thx for the trouble. :)